A SPECTROSCOPIC STUDY OF CaOCH_{3} USING THE PUMP/PROBE MICROWAVE AND THE MOLEC-ULAR BEAM/OPTICAL STARK TECHNIQUES

\underline{KEI-ICHI NAMIKI}, J. SCOTT ROBINSON, AND TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287.

The Stark effect on the ^{q}R_{22}(0,0.5) (ν= 17682.9251 cm^{−1}) and ^{q}P_{11}(0,1.5) (ν= 17682.1966 cm^{−1}) branch features of the (0,0)B^{2}A_{1} − X^{2}A_{1} band system of calcium methoxide, CaOCH_{3}, was measured and analyzed to give the permanent electronic dipole moments, μ, of 1.58(8)D and 1.21(5)D for the X^{2}A_{1} and B^{2}A_{1} states, respectively. The dipole moments are compared with other monovalent calcium compounds and those predicted from a simple electrostatic model. Pure rotational transitions in the X^{2}A_{1} state were recorded using the pump/probe microwave-optical double resonance technique. The proton magnetic hyperfine splitting pattern confirms that the symmetry of the ground electronic state is C_{3v}. The determined small negative value for the Fermi contact parameter (a_{F} = -0.419 MHz) is interpreted in terms of spin polarization effects. The determined spin-rotational parameter ((ϵ_{bb} + ϵ_{cc})/2 = 12.45 MHz ) is compared to that of other monovalent calcium compounds and interpreted in terms of the proposed state distribution.